OpenStudy (anonymous):

WILL FAN AND MEDAL! What is the solution to the system of equations? x+y=5 x-2y=2 (_,_)

jabez177 (jabez177):

I'm sorry. I'm not feeling my Math vibe at the moment. @just_one_last_goodbye

OpenStudy (anonymous):

That's okay, thanks though.

OpenStudy (anonymous):

@TheSmartOne could you please help me again? I didn't quite understand, do I subtract 2 from 5?

OpenStudy (anonymous):

I'm not very good with these types of questions at all, sorry

TheSmartOne (thesmartone):

x - x + y - 2y = 5 - 2 Add the like terms.

OpenStudy (anonymous):

so -2 and 2? sorry

TheSmartOne (thesmartone):

What is y - 2y = . y - 2y = y (1 -2) = ?

OpenStudy (anonymous):

-1?

TheSmartOne (thesmartone):

It would be y (-1) = -1y = -y

TheSmartOne (thesmartone):

What is 5 - 2 = ?